OpenStudy (anonymous):

cosx-cosx/1-tanx = sinxcosx/sinx-cosx halp :c

OpenStudy (anonymous):

cos X = adj / hyp = b / c , sec X = hyp / adj = c / b ,

OpenStudy (anonymous):

need to prove it :/

OpenStudy (anonymous):

cscX = 1 / sinX sinX = 1 / cscX secX = 1 / cosX cosX = 1 / secX tanX = 1 / cotX cotX = 1 / tanX tanX = sinX / cosX cotX = cosX / sinX

OpenStudy (anonymous):

cosx - cosx ?